OpenStudy (anonymous):

Solve for x on [0.2π)

geerky42 (geerky42):

more information needed

OpenStudy (anonymous):

solve for x on [0,2π) sin^2 x=sinx

geerky42 (geerky42):

Try to use this identity: \(\sin^2x=\dfrac{1-\cos(2x)}{2}\)

OpenStudy (anonymous):

why?

OpenStudy (aaronq):

if you use that identity you can simplify the equation so you can solve for x.

OpenStudy (xapproachesinfinity):

or do this sin^2x-sinx=0 ==> sinx(sinx-1)=0 now solve for x